OpenStudy (rachie19):

the atomic radius of a carbon atom is 7.72* 10-2 nm express length in micrometers

OpenStudy (cuanchi):

10^9 nm = 10^6 um

OpenStudy (cuanchi):

you can made a conversion factor and multiply your value by the conversion factor 10^9 nm/ 10^6 um = 10^6um /10^9 nm 7.72 x 10^(-2) nm x 10^6um /10^9 nm = ???

OpenStudy (rachie19):

ok i got it thank you =)